OpenStudy (anonymous):

Part 1: Using complete sentences, explain what the discriminant is and what it tells you about the solutions of a quadratic equation. (4 points) Part 2: Provide a unique example to back up your explanation. (4 points)

hero (hero):

Are you taking a test?

OpenStudy (anonymous):

The general form of a quadratic equation is ax^2 + bx + c = 0 where a, b and c are numbers. There is a formula for the roots and that is( -b+- sqrt(b^2-4ac) / 2a The disvriminant is b^2-4ac. If discriminant is > 0 there are two real roots, if equal to 0 there are two equal real roots and if it is < 0 then the roots are imaginary. eg solve x^2 + 5x + 13 = 0 Discriminant = 5^2 - 4*1* 13 = -27 the sqrt of -27 is imaginary no real roots